Use Laplace Transform to solve the given initial-value problem.

y''' − 16y' = e^t

y(0) = 0

y''(0) = 0

y'(0) = 4
Answer #1

8 form to both side of the given Since h(y) = sh(Y) - Y (o) then h (yi) = sh(y) - Y(o) >> A[y) = ($1LY)-yo)) - YO) 7 h (4)

ਹੈ (. |6 ਜੀ +47 ਕੇ LL get (5) + ਈ 15 16 ) 4Y > Lly) Vib (1) = £3-1 581) ੫ S4॥ S1 - ਨੂੰ 6 ਹੈ | ਪy = $ ਛਾਏ ਨੂੰ ਸੀ - + ਤੇ - ਹੈ

ਜLCy) = ਦਰ ਚ + ਏ – ਹਰਿ ਹੈ ! . 210 ਉਤਰ ) - $1 ਦੀ * Y = cash 4t + A sinh ft-t et tet L La) = cash at Ill standa á sinh at